OpenStudy (anonymous):

The equation sin ( (PI/3) - x ) is equal to ____. A. 1/2( sqrt(3) cosx + sinx) B. 1/2(sqrt(3) cosx - sinx) C. 1/2(cosx- sqrt(3) sinx) D. 1/2 (cosx + sqrt(3) sinx)

zepdrix (zepdrix):

The Angle Difference Formula for Sine tells us that... \[\large \sin(\alpha-\beta)=\sin \alpha \cos \beta- \sin \beta \cos \alpha\]

zepdrix (zepdrix):

Applying this to our problem gives us... \[\large \sin(\frac{ \pi }{ 3 }-x)=\sin\frac{ \pi }{ 3 }\cos x- \sin x \cos\frac{ \pi }{ 3 }\]

zepdrix (zepdrix):

From there, you'll need t remember your special angles.. One of them is 1/2, the other is going to be sqrt3/2, you'll have to figure out which is which. :D After you fill in the pieces, you'll be able to factor 1/2 out of each term, and it should match one of your answers. Yay!

OpenStudy (anonymous):

...you say yay ..im still like wat? ..

zepdrix (zepdrix):

lolol, confused on what part? :O

OpenStudy (anonymous):

all of it ._.

zepdrix (zepdrix):

We started with a problem that contains a weird looking angle (The thing inside the sine function is an angle, we apply trig functions to angles). We have a formula that allows us to mess with the angle, split it up into pieces :O Hmmm trig is a little tricky to explain c:

zepdrix (zepdrix):

In the formula that I stated earlier, in our problem the corresponding pieces are \[\huge \alpha = \frac{ \pi }{ 3 }, \beta= x\]

OpenStudy (anonymous):

..yeah it is :/ um what?

zepdrix (zepdrix):

 You've learned these formulas yes? :O Just confused on how to apply them?